OpenStudy (anonymous):

a student signs up for 3 courses for this semester. the probability of getting an a in class 1 is 0.3. the probability of getting an a in class 2 is 0.4. the probability of getting an a in class 3 is 0.5.

OpenStudy (anonymous):

a) what is the probability of getting 3 As this semester? b) what is the probability of getting 2 As this semester? c) what is the probability of getting at least 1 A this semester?

OpenStudy (anonymous):

a) multiply the three numbers together \(.3\times .4\times .5\) assuming independence

OpenStudy (anonymous):

yeah i was thinking the same thing for A but what about the others

OpenStudy (anonymous):

b) two A's you have to work by cases

OpenStudy (anonymous):

class 1, class 2, class 3 A A not A \(.3\times .4\times .5\) A not A A \(.3\times .6\times .5\) not A A A \(.7\times .4\times .5\) find these three numbers, add them up

OpenStudy (anonymous):

c) at least on A is easiest found by computing the probability that you get exactly no A and subtract that number from 1

OpenStudy (anonymous):

*one A

OpenStudy (anonymous):

i.e. \(1-.7\times .6\times .5\)

OpenStudy (anonymous):

Thanks big time!